A cylindrical disk of wood weighing πŸ“πŸŽ.𝟎 [N] and having a diameter of πŸπŸ“.𝟎 [cm] floats on a liquid of density 𝝆=𝟎.πŸ•πŸ“πŸŽ [g/cm3]. The cylinder of liquid is πŸ”πŸŽ.𝟎 [cm] high and has a diameter that is the same as that of the wood. What is the gauge pressure halfway down the liquid?


1
Expert's answer
2021-06-01T10:50:58-0400

The gauge pressure is the pressure ignoring the atmospheric pressure. So, the wooden cylinder creates a pressure of


"p_1=\frac FA=\frac{4F}{\pi d^2}."

The pressure of liquid at the depth of 30 cm is


"p_2=\rho gh."

The gauge pressure:


"P=p_1+p_2=\rho gh+\frac{4F}{\pi d^2}=3223.6\text{ Pa}."